# RetCompletion:High-Speed Inference Image Completion with Retentive Network

### v1

@@ Abstract
Time cost is a major challenge in achieving high-quality pluralistic image completion. Recently, the Retentive Network (RetNet) in natural language processing offers a novel approach to this problem with its low-cost inference capabilities. Inspired by this, we apply RetNet to the pluralistic image completion task in computer vision. We present RetCompletion, a two-stage framework. In the first stage, we introduce Bi-RetNet, a bidirectional sequence information fusion model that integrates contextual information from images. During inference, we employ a unidirectional pixel-wise update strategy to restore consistent image structures, achieving both high reconstruction quality and fast inference speed. In the second stage, we use a CNN for low-resolution upsampling to enhance texture details. Experiments on ImageNet and CelebA-HQ demonstrate that our inference speed is 10$\times$ faster than ICT and 15$\times$ faster than RePaint. The proposed RetCompletion significantly improves inference speed and delivers strong performance, especially when masks cover large areas of the image.
1 Introduction
Pluralistic image completion, also known as image inpainting, is a crucial research area with various applications, including object removal and photo restoration Barnes et al. (2009); Criminisi et al. (2004); Dale et al. (2009); Wan et al. (2020). CNN-based methods liu2021; Iizuka et al. (2017); Li et al. (2017) have demonstrated impressive results by capturing local texture patterns, but they often struggle to model global structures, leading to suboptimal image reconstruction quality. To overcome this limitation, researchers have introduced hybrid models combining Transformers and CNNs Wan et al. (2021); Zheng et al. (2022); Li et al. (2022). While these approaches significantly improve reconstruction quality and produce diverse results by modeling the underlying data distribution, Transformer-based pixel-wise generation involves extensive feature fusion calculations. This computational overhead increases inference time, limiting the practicality of these methods, especially in real-time applications. Therefore, developing algorithms that maintain high-quality reconstruction while improving computational efficiency remains a critical challenge in this domain.
Recently, Retentive Network (RetNet) Sun et al. (2023) has shown substantial potential in natural language processing due to its multi-scale retention mechanism, which bridges parallel training and recurrent inference. This capability enables RetNet to process information efficiently, even in pixel-wise generation tasks. However, applying RetNet directly to vision tasks presents challenges, as image information is not unidirectional like language data.
@@ 2 Related Work
The significance of Pluralistic Image Completion lies in providing a diverse approach to image processing, allowing for the creation of images with different styles and effects, enriching the toolbox in creative and design fields, and supporting diverse choices in decision-making processes. PIC Zheng et al. (2019) employs a dual-path framework based on probabilistic principles: one is the reconstructive path, which utilizes the given ground truth to obtain prior information about the missing parts and reconstructs the original image from this distribution. The other is the generative path, where the conditional prior is coupled with the distribution from the reconstructive path. ICT Wan et al. (2021) directly optimizes the log-likelihood in the discrete space in the first transformer-based stage without the need for additional assumptions. RePaint Lugmayr et al. (2022) applies the Diffusion model to the image inpainting task, using a pre-trained unconditional DDPM Ho et al. (2020) as the generative prior and modifying the reverse diffusion iterations by sampling the unmasked regions from the given image information. Since this technique doesn’t alter or condition the original DDPM Ho et al. (2020) network itself, the model can generate high-quality and diverse output images for any inpainting scenario.
@@
Retentive Network Sun et al. (2023) introduces the retention mechanism with a dual form of recurrence and parallelism. It has three computation paradigms,i.e., parallel, recurrent, and chunkwise recurrent. We can train parallelly by using parallel paradigm while conducting inference recurrently using recurrent and chunkwise paradigms. The retention mechanism utilizes a rotation-based positional encoding along with a decay term to effectively model the position information of tokens, known as xPos Sun et al. (2022), a relative position embedding proposed for Transformer. We attempt to extend this method to two-dimensional images.
3 Methods
@@
3.1 Retentive Network
Retentive Network (RetNet) is a powerful architecture initially designed for natural language processing, which combines parallel and recurrent representations to efficiently handle sequential data. Its key advantage lies in its ability to balance parallel training and recurrent inference, enabling fast computation even in complex tasks.
RetNet models sequences in a recurrent manner, where the hidden state at each step is computed as:
$s_{n}=As_{n-1}+K_{n}^{\top}v_{n}$ (1)
This allows RetNet to accumulate information over time while maintaining efficient updates.
Additionally, RetNet can be diagonalized to simplify the recurrence into a more efficient form using positional encoding, further improving its speed and accuracy:
$o_{n}=\sum_{m=1}^{n}\gamma^{n-m}(Q_{n}e^{in\theta})(K_{m}e^{im\theta})^{\dagger}v_{m}$ (2)
The RetNet framework incorporates three key representations:
Parallel Representation
Parallelization allows RetNet to achieve linear complexity during training, greatly speeding up the process. This is computed as:
$Retention(X)=(QK^{T}\odot D)V$ (3)
Recurrent Representation
During inference, RetNet uses its recurrent form, which achieves O(1) complexity, enabling fast pixel-wise inference:
$Retention(X_{n})=Q_{n}S_{n}$ (4)
Chunkwise Recurrent Representation
This combines both parallel and recurrent approaches, allowing for accelerated computations in large-scale tasks:
$Retention(X_{[i]})=(Q_{[i]}K_{[i]}^{T}\odot D)V_{[i]}+Q_{[i]}R_{[i-1]}$ (5)
This unique combination of representations allows RetNet to efficiently process large datasets, making it particularly well-suited for pixel-wise image completion tasks, where both speed and accuracy are critical.
3.2 Preprocessing
@@
RGB color channels typically exist in a high-dimensional space ($256^{3}$ colors), which makes direct processing computationally expensive. To handle this, we generate a compact visual vocabulary by applying K-Means clustering on the entire ImageNet dataset Russakovsky et al. (2015), reducing the space to 512 representative colors. For each unmasked pixel, we map its color to the nearest representative color from this vocabulary. The image is then raster-scanned and reshaped into a sequence, which is necessary for the RetNet model:
@@
3.3 Appearance Priors Reconstruction by Bi-RetNet
@@
Similar to BERT Devlin et al. (2018), we employ the Masked Language Model (MLM) objective to optimize the RetNet. The loss function minimizes the negative log-likelihood of the masked pixels, ensuring that the model learns to predict missing regions accurately:
@@ 3.4 Parallel Training
3.5 Pixel-wise Inference
@@
The inference process is detailed in Algorithm 1, where the model integrates the forward and backward information for each pixel prediction and updates the RetNet’s retention state after each step:
@@
This pixel-wise inference strategy allows our model to efficiently update and refine each pixel prediction, improving both speed and accuracy compared to Transformer-based methods.
3.6 Guided Upsampling
After reconstructing the appearance priors, we reshape the sequence $X\in\mathbb{R}^{L^{2}\times 3}$ into $I_{t}\in\mathbb{R}^{L\times L\times 3}$, which represents a low-resolution image. We then upscale this image to the original resolution $H\times W\times 3$. Following ICT Wan et al. (2021), we employ a CNN-based guided upsampling network, as CNNs have shown excellent performance in texture reconstruction. The upsampling network is composed of an encoder, decoder, and residual blocks. First, we upsample $I_{t}$ to the original resolution using bilinear interpolation, and then we feed the upsampled image along with the original image and mask into the upsampling network as:
$I_{\text{pred}}=\mathit{F}_{\delta}(I^{\uparrow}_{t}\smallfrown I_{m})\in\mathbb{R}^{H\times W\times 3}$ (16)
where $\mathit{F}$ represents the upsampling network with parameters $\delta$.
We apply both $L_{1}$ loss between $I_{\text{pred}}$ and $I$, and adversarial loss to train the upsampling network as:
$L_{L_{1}}=\mathbb{E}[|I_{\text{pred}}-I|_{1}]$ (17)
$L_{\text{adv}}=\mathbb{E}[\log(1-\mathit{D}\omega(I{\text{pred}}))]+\mathbb{E}[\log\mathit{D}_{\omega}(I)]$ (18)
where $\mathit{D}$ is the discriminator with parameters $\omega$.
The upsampling network $\mathit{F}$ and discriminator $\mathit{D}$ are trained with the following optimization objective:
$\min_{\mathit{F}}\max_{\mathit{D}}L_{\text{upsample}}(\delta,\omega)=\alpha_{1}L_{L_{1}}+\alpha_{2}L_{\text{adv}}$ (19)
4 Experiments
In our experiments, we evaluate the performance of the proposed method using two datasets: CelebA-HQ Karras et al. (2017) and ImageNet Russakovsky et al. (2015). We perform both quantitative and qualitative evaluations to assess the quality of image completion and the inference speed. Quantitative comparisons are conducted against other state-of-the-art methods in terms of image quality and computational efficiency, while qualitative comparisons are based on user feedback. Note that all qualitative and quantitative results reported in this paper are based on a fixed image resolution of 256 pixels.
4.1 Implementation Details
To ensure fair comparisons across different datasets and methods, we follow the same configuration as ICT Wan et al. (2021) for all experiments, as shown in Table 1. For the CelebA-HQ Karras et al. (2017) and ImageNet Russakovsky et al. (2015) datasets, we retain the original training and test splits. Additionally, we employ PConv Liu et al. (2018) to generate diverse masks during training to simulate various image occlusion scenarios.
@@ 4.1 Implementation Details
CelebA-HQ Karras et al. (2017) 8 512 30 48$\times$48
ImageNet Russakovsky et al. (2015) 8 1024 35 32$\times$32
Table 1: Retention Network parameter setting across different experiment. h: Head number. d: The dimension of embedding space. N: Number of retention layers. L: The length of appearance prior.
Dataset CelebA-HQ Karras et al. (2019) ImageNetRussakovsky et al. (2015)
@@
PIC Zheng et al. (2019) Wide 23.781 0.883 0.164 23.765 0.819 0.234
LaMa Suvorov et al. (2022) 27.581 0.928 0.045 26.099 0.865 0.105
RePaint Lugmayr et al. (2022) 27.496 0.931 0.059 25.768 0.859 0.134
$ICT^{*}$ Wan et al. (2021) 26.897 0.922 0.069 25.545 0.848 0.125
ICT Wan et al. (2021) 27.139 0.932 0.063 25.886 0.862 0.107
@@
PIC Zheng et al. (2019) Narrow 25.823 0.901 0.062 24.091 0.823 0.098
LaMa Suvorov et al. (2022) 28.684 0.942 0.028 26.892 0.902 0.061
RePaint Lugmayr et al. (2022) 28.547 0.938 0.028 26.908 0.906 0.064
$ICT^{*}$ Wan et al. (2021) 28.242 0.932 0.041 26.887 0.898 0.079
ICT Wan et al. (2021) 28.551 0.944 0.036 26.902 0.903 0.073
@@
PIC Zheng et al. (2019) Half 21.484 0.852 0.238 19.498 0.708 0.354
LaMa Suvorov et al. (2022) 25.208 0.905 0.138 23.513 0.756 0.254
RePaint Lugmayr et al. (2022) 24.846 0.902 0.165 23.498 0.762 0.304
$ICT^{*}$ Wan et al. (2021) 24.356 0.896 0.179 23.476 0.748 0.278
ICT Wan et al. (2021) 24.798 0.906 0.166 23.502 0.753 0.255
@@
PIC Zheng et al. (2019) Center 25.580 0.887 0.153 23.806 0.816 0.167
LaMa Suvorov et al. (2022) 28.529 0.940 0.039 26.276 0.886 0.086
RePaint Lugmayr et al. (2022) 28.556 0.940 0.041 26.304 0.886 0.093
$ICT^{*}$ Wan et al. (2021) 28.409 0.935 0.058 26.198 0.879 0.103
ICT Wan et al. (2021) 28.496 0.942 0.052 26.282 0.888 0.096
@@
PIC Zheng et al. (2019) Expand 18.893 0.798 0.576 17.364 0.652 0.712
LaMa Suvorov et al. (2022) 23.382 0.878 0.342 20.384 0.697 0.534
RePaint Lugmayr et al. (2022) 23.376 0.882 0.435 20.439 0.702 0.629
$ICT^{*}$ Wan et al. (2021) 23.298 0.876 0.446 20.126 0.683 0.562
ICT Wan et al. (2021) 23.379 0.879 0.432 20.324 0.698 0.544
@@
Table 2: Quantitative results on CelebA-HQ Karras et al. (2017) and ImageNet Russakovsky et al. (2015) datasets with different mask types. All the pluralistic image completion methods use Top-1 sampling. The models with * indicate the prediction method that uses all pixel points simultaneously, while the models without * indicate the prediction method that uses pixel-by-pixel prediction.
4.2 Quantitative Comparisons
@@
We quantitatively compare our method against several state-of-the-art (SOTA) image completion techniques using peak signal-to-noise ratio (PSNR), structural similarity index (SSIM), and learned perceptual image patch similarity (LPIPS). Experiments are conducted on CelebA-HQ Karras et al. (2017) and ImageNet Russakovsky et al. (2015) datasets with five distinct mask types to assess performance across different occlusion patterns. For all pluralistic image completion methods, Top-1 sampling is applied during testing.
@@
The results of the quantitative experiments are presented in Table 2. It is clear that our method consistently outperforms most existing SOTA methods across various mask types and datasets. A notable observation is the significant difference between pixel-wise inference and simultaneous pixel estimation (denoted with * in the table). Our pixel-wise inference approach demonstrates clear superiority in terms of both image quality and perceptual similarity, as evidenced by the improvements in PSNR, SSIM, and LPIPS metrics. This emphasizes the effectiveness of our method, particularly in scenarios with complex occlusions.
While this section highlights the quality improvements, the efficiency of our method is also noteworthy. Our pixel-wise inference strategy not only yields better image reconstruction but also achieves faster inference times compared to methods that predict all pixels simultaneously. We will explore this speed advantage in more detail in the subsequent section, where we provide a visual comparison of the inference time between our method and others.
Mask types
We utilize five distinct mask types in our experiments. The Wide and Narrow masks are adapted from LaMa Suvorov et al. (2022), representing occlusions of different widths across the image. The Half mask randomly occludes either the top, bottom, left, or right part of the image. The Center mask covers a central $64\times 64$ region, while the Expand mask occludes all regions except for the central part, covering the majority of the image.
Figure 2: Sample images for user study.
4.3 User study
@@
To enhance the assessment of subjective quality, we additionally perform a user study to compare our method against other baseline approaches. We randomly select 50 images and apply various masks to each. Employing different image completion methods, including pluralistic image completion methods, we consistently used the Top-1 sampling result. Specifically, we present a set of five images generated by MED Liu et al. (2020), PIC Zheng et al. (2019), EC Nazeri et al. (2019), ICT Wan et al. (2021), and our method for each image. Users are then asked to rank the top three images that appear most realistic. Finally, we calculate the percentage of times each method ranked within the top three. Sample images for user study are shown in Figure.  2.
@@
The results obtained from 200 users are shown in Figure 3(a). The results show that our method significantly outperforms the PIC method in terms of visual quality. Additionally, our method shows a slight advantage over the ICTWan et al. (2021), LaMa Suvorov et al. (2022), and RePaintLugmayr et al. (2022) methods. This demonstrates the superiority of our method in terms of visual perception.
@@
4.4 Inference Time
To ensure a fair comparison, we measured the pixel-wise completion inference time of ICT Wan et al. (2021), RePaint Lugmayr et al. (2022), and our proposed method on the ImageNet Russakovsky et al. (2015) dataset using a GeForce RTX 4090 GPU. The results, shown in Figure 3(b), demonstrate that our method achieves significantly lower inference times compared to the other two methods, particularly at higher mask rates.
For the RePaint method, although its inference time remains consistent regardless of the mask rate, it requires numerous iterative denoising steps, leading to a substantial overall inference time. On the other hand, ICT Wan et al. (2021) recalculates attention for each pixel estimation, causing the inference time to increase linearly as the mask rate rises. In contrast, our method leverages a recurrent computation paradigm, where only the information of changed pixels is updated and integrated into the state $S$. As a result, our method shows only a minimal increase in inference time across different mask rates, ensuring consistently high-speed performance.
5 Limitations
As illustrated in Figure  2, current image completion methods encounter significant difficulties when dealing with more challenging mask types. For example, the performance on Expand-type masks is particularly subpar, especially when applied to highly diverse datasets such as ImageNet. This highlights a notable gap in the effectiveness of existing approaches. Recognizing this limitation, our future research will place a stronger emphasis on improving the robustness and accuracy of image completion techniques for these complex mask scenarios.
6 Conclusion

### v2

@@ Abstract
Time cost is a major challenge in achieving high-quality pluralistic image completion. Recently, the Retentive Network (RetNet) in natural language processing offers a novel approach to this problem with its low-cost inference capabilities. Inspired by this, we apply RetNet to the pluralistic image completion task in computer vision. We present RetCompletion, a two-stage framework. In the first stage, we introduce Bi-RetNet, a bidirectional sequence information fusion model that integrates contextual information from images. During inference, we employ a unidirectional pixel-wise update strategy to restore consistent image structures, achieving both high reconstruction quality and fast inference speed. In the second stage, we use a CNN for low-resolution upsampling to enhance texture details. Experiments on ImageNet and CelebA-HQ demonstrate that our inference speed is 10$\times$ faster than ICT and 15$\times$ faster than RePaint. The proposed RetCompletion significantly improves inference speed and delivers strong performance.
Index Terms:
Pluralistic Image Completion, Retentive Network, Fast Inference, Bi-RetNet
I Introduction
Pluralistic image completion, also known as image inpainting, is a crucial research area with various applications, including object removal and photo restoration [1, 2, 3, 4]. CNN-based methods [5] have demonstrated impressive results by capturing local texture patterns, but they often struggle to model global structures, leading to suboptimal image reconstruction quality. To overcome this limitation, researchers have introduced hybrid models combining Transformers and CNNs [6, 7, 8]. While these approaches significantly improve reconstruction quality and produce diverse results by modeling the underlying data distribution, Transformer-based pixel-wise generation involves extensive feature fusion calculations. This computational overhead increases inference time, limiting the practicality of these methods, especially in real-time applications. Therefore, developing algorithms that maintain high-quality reconstruction while improving computational efficiency remains a critical challenge in this domain.
Recently, Retentive Network (RetNet) [9] has shown substantial potential in natural language processing due to its multi-scale retention mechanism, which bridges parallel training and recurrent inference. This capability enables RetNet to process information efficiently, even in pixel-wise generation tasks. However, applying RetNet directly to vision tasks presents challenges, as image information is not unidirectional like language data.
@@ II Related Work
The significance of Pluralistic Image Completion lies in providing a diverse approach to image processing, allowing for the creation of images with different styles and effects, enriching the toolbox in creative and design fields, and supporting diverse choices in decision-making processes. PIC [10] employs a dual-path framework based on probabilistic principles: one is the reconstructive path, which utilizes the given ground truth to obtain prior information about the missing parts and reconstructs the original image from this distribution. The other is the generative path, where the conditional prior is coupled with the distribution from the reconstructive path. ICT [6] directly optimizes the log-likelihood in the discrete space in the first transformer-based stage without the need for additional assumptions. RePaint [11] applies the Diffusion model to the image inpainting task, using a pre-trained unconditional DDPM [12] as the generative prior and modifying the reverse diffusion iterations by sampling the unmasked regions from the given image information. Since this technique doesn’t alter or condition the original DDPM [12] network itself, the model can generate high-quality and diverse output images for any inpainting scenario.
@@
Retentive Network [9] introduces the retention mechanism with a dual form of recurrence and parallelism. It has three computation paradigms,i.e., parallel, recurrent, and chunkwise recurrent. We can train parallelly by using parallel paradigm while conducting inference recurrently using recurrent and chunkwise paradigms. The retention mechanism utilizes a rotation-based positional encoding along with a decay term to effectively model the position information of tokens, known as xPos [13], a relative position embedding proposed for Transformer. We attempt to extend this method to two-dimensional images.
III Methods
@@
III-A Preprocessing
@@
RGB color channels typically exist in a high-dimensional space ($256^{3}$ colors), which makes direct processing computationally expensive. To handle this, we generate a compact visual vocabulary by applying K-Means clustering on the entire ImageNet dataset [14], reducing the space to 512 representative colors. For each unmasked pixel, we map its color to the nearest representative color from this vocabulary. The image is then raster-scanned and reshaped into a sequence, which is necessary for the RetNet model:
@@
III-B Appearance Priors Reconstruction by Bi-RetNet
@@
Similar to BERT [15], we employ the Masked Language Model (MLM) objective to optimize the RetNet. The loss function minimizes the negative log-likelihood of the masked pixels, ensuring that the model learns to predict missing regions accurately:
@@ III-C Parallel Training
III-D Pixel-wise Inference
@@
The inference process is detailed in Algorithm 1, where the model integrates the forward and backward information for each pixel prediction and updates the RetNet’s retention state after each step.
@@
III-E Guided Upsampling
ollowing ICT [6], we use a CNN-based guided upsampling network to reconstruct the high-resolution image. After reshaping the appearance priors into a low-resolution image, we first apply bilinear interpolation to upscale it to the original resolution. The upsampled image, along with the original image and mask, is then fed into the upsampling network. We train the network using both $L_{1}$ loss and adversarial loss, with the upsampling network and discriminator being optimized jointly.
IV Experiments
Figure 2: Sample images for user study.
In our experiments, we evaluate the performance of the proposed method using two datasets: CelebA-HQ [16] and ImageNet [14]. We perform both quantitative and qualitative evaluations to assess the quality of image completion and the inference speed. Quantitative comparisons are conducted against other state-of-the-art methods in terms of image quality and computational efficiency, while qualitative comparisons are based on user feedback. Note that all qualitative and quantitative results reported in this paper are based on a fixed image resolution of 256 pixels.
IV-A Implementation Details
To ensure fair comparisons across different datasets and methods, we follow the same configuration as ICT [6] for all experiments, as shown in Table I. For the CelebA-HQ [16] and ImageNet [14] datasets, we retain the original training and test splits. Additionally, we employ PConv [17] to generate diverse masks during training to simulate various image occlusion scenarios.
@@ IV-A Implementation Details
CelebA-HQ [16] 8 512 30 48$\times$48
ImageNet [14] 8 1024 35 32$\times$32
TABLE I: Retention Network parameter setting across different experiment. h: Head number. d: The dimension of embedding space. N: Number of retention layers. L: The length of appearance prior.
Dataset CelebA-HQ [18] ImageNet[14]
@@
PIC [10] Wide 23.781 0.883 0.164 23.765 0.819 0.234
LaMa [19] 27.581 0.928 0.045 26.099 0.865 0.105
RePaint [11] 27.496 0.931 0.059 25.768 0.859 0.134
$ICT^{*}$ [6] 26.897 0.922 0.069 25.545 0.848 0.125
ICT [6] 27.139 0.932 0.063 25.886 0.862 0.107
@@
PIC [10] Narrow 25.823 0.901 0.062 24.091 0.823 0.098
LaMa [19] 28.684 0.942 0.028 26.892 0.902 0.061
RePaint [11] 28.547 0.938 0.028 26.908 0.906 0.064
$ICT^{*}$ [6] 28.242 0.932 0.041 26.887 0.898 0.079
ICT [6] 28.551 0.944 0.036 26.902 0.903 0.073
@@
PIC [10] Half 21.484 0.852 0.238 19.498 0.708 0.354
LaMa [19] 25.208 0.905 0.138 23.513 0.756 0.254
RePaint [11] 24.846 0.902 0.165 23.498 0.762 0.304
$ICT^{*}$ [6] 24.356 0.896 0.179 23.476 0.748 0.278
ICT [6] 24.798 0.906 0.166 23.502 0.753 0.255
@@
PIC [10] Center 25.580 0.887 0.153 23.806 0.816 0.167
LaMa [19] 28.529 0.940 0.039 26.276 0.886 0.086
RePaint [11] 28.556 0.940 0.041 26.304 0.886 0.093
$ICT^{*}$ [6] 28.409 0.935 0.058 26.198 0.879 0.103
ICT [6] 28.496 0.942 0.052 26.282 0.888 0.096
@@
PIC [10] Expand 18.893 0.798 0.576 17.364 0.652 0.712
LaMa [19] 23.382 0.878 0.342 20.384 0.697 0.534
RePaint [11] 23.376 0.882 0.435 20.439 0.702 0.629
$ICT^{*}$ [6] 23.298 0.876 0.446 20.126 0.683 0.562
ICT [6] 23.379 0.879 0.432 20.324 0.698 0.544
@@
TABLE II: Quantitative results on CelebA-HQ [16] and ImageNet [14] datasets with different mask types. All the pluralistic image completion methods use Top-1 sampling. The models with * indicate the prediction method that uses all pixel points simultaneously, while the models without * indicate the prediction method that uses pixel-by-pixel prediction.
IV-B Quantitative Comparisons
@@
We quantitatively compare our method against several state-of-the-art (SOTA) image completion techniques using peak signal-to-noise ratio (PSNR), structural similarity index (SSIM), and learned perceptual image patch similarity (LPIPS). Experiments are conducted on CelebA-HQ [16] and ImageNet [14] datasets with five distinct mask types to assess performance across different occlusion patterns. For all pluralistic image completion methods, Top-1 sampling is applied during testing.
@@
The results of the quantitative experiments are presented in Table II, which shows performance for the Wide and Expand mask types. Our method consistently outperforms most existing SOTA methods across various mask types and datasets, demonstrating clear superiority in terms of image quality and perceptual similarity, as evidenced by improvements in PSNR, SSIM, and LPIPS metrics. A notable observation is the significant difference between pixel-wise inference and simultaneous pixel estimation (denoted with * in the table). Our pixel-wise inference approach performs better, especially in scenarios with complex occlusions. We employ five distinct mask types in our experiments: Wide and Narrow masks (adapted from LaMa [19]), Half mask (random occlusion of parts of the image), Center mask (covering a central $64\times 64$ region), and Expand mask (occluding everything except the central region). For the complete results across all mask types, please refer to Appendix.
IV-C User study
@@
To enhance the assessment of subjective quality, we additionally perform a user study to compare our method against other baseline approaches. We randomly select 50 images and apply various masks to each. Employing different image completion methods, including pluralistic image completion methods, we consistently used the Top-1 sampling result. Specifically, we present a set of five images generated by MED [20], PIC [10], EC [21], ICT [6], and our method for each image. Users are then asked to rank the top three images that appear most realistic. Finally, we calculate the percentage of times each method ranked within the top three. Sample images for user study are shown in Figure.  2.
@@
The results obtained from 200 users are shown in Figure 3(a). The results show that our method significantly outperforms the PIC method in terms of visual quality. Additionally, our method shows a slight advantage over the ICT[6], LaMa [19], and RePaint[11] methods. This demonstrates the superiority of our method in terms of visual perception.
@@
IV-D Inference Time
To ensure a fair comparison, we measured the pixel-wise completion inference time of ICT [6], RePaint [11], and our method on the ImageNet [14] dataset using a GeForce RTX 4090 GPU. The results, shown in Figure 3(b), indicate that our method achieves significantly lower inference times, especially at higher mask rates.
RePaint maintains consistent inference time but requires numerous denoising iterations, leading to a higher overall time. ICT [6] recalculates attention for each pixel, causing a linear increase in time as the mask rate rises. In contrast, our method uses a recurrent computation paradigm, where only changed pixels are updated, resulting in minimal increases in inference time and consistently high-speed performance.
V Conclusion